# Revisiting Weight Averaging for Model Merging

### v1

@@ Abstract
Model merging aims to build a multi-task learner by combining the parameters of individually fine-tuned models without additional training. While a straightforward approach is to average model parameters across tasks, this often results in suboptimal performance due to interference among parameters across tasks. In this paper, we present intriguing results that weight averaging implicitly induces task vectors centered around the weight averaging itself and that applying a low-rank approximation to these centered task vectors significantly improves merging performance. Our analysis shows that centering the task vectors effectively separates core task-specific knowledge and nuisance noise within the fine-tuned parameters into the top and lower singular vectors, respectively, allowing us to reduce inter-task interference through its low-rank approximation. We evaluate our method on eight image classification tasks, demonstrating that it outperforms prior methods by a significant margin, narrowing the performance gap with traditional multi-task learning to within 1-3%.
@@ 1 Introduction
Model merging has emerged as an efficient way for constructing multi-task learners [18]. Unlike traditional multi-task learning approaches that directly train a single model on multiple tasks [1], model merging leverages individually fine-tuned models and fuses their parameters to create a model that preserves their original capabilities. This approach eliminates the need to prepare training data or store separate sets of parameters for each task, thereby reducing the costs associated with modern deep neural networks, which often require large amounts of data and numerous parameters. Consequently, model merging has been favored in various applications, including federated learning [25], model compression [32], and continual learning [10].
@@
To mitigate the limitations of the weight averaging, recent approaches have leveraged the task arithmetic framework [12], which allows for the extrapolation of the parameters. In this framework, models are assumed to be fine-tuned from a common initialization, enabling the definition of task vectors as the directions pointing from the initialization to the fine-tuned parameters in the parameter space. Among several arithmetic operations on task vectors, addition with a scaling coefficient has shown to be effective in merging models trained on various tasks. Subsequent approaches have improved this arithmetic by either resolving interference between task vectors [35] or applying test-time adaptation techniques for better scaling coefficient [36].
In this paper, we revisit the weight averaging strategy through the lens of task arithmetic. We begin by formulating weight averaging as a task arithmetic that induces centered task vectors around the weight average itself. We then observe that applying low-rank approximations on the centered task vectors dramatically improves the performance of the merged model produced by the corresponding task arithmetic. When an appropriate rank is chosen, this approach largely outperforms the original task arithmetic and even advanced task vector-based variants. We also note that the optimal rank can be consistently found throughout different experimental settings. To understand this surprisingly high performance, we provide both theoretical insights and empirical evidence based on the spectral analysis of task vectors. We conjecture that centering the task vectors leads to the isolation of core task-specific knowledge from the interfering noise present in their top and lower singular vectors.
Based on this observation, we propose a novel training-free model merging approach called Centered Arithmetic with Rank-reduced Task vectors (CART), which consistently outperforms existing baselines across various tasks and model architectures. This method can be seamlessly integrated into existing merging approaches based on task vectors, such as those using test-time adaptation. In our experiments, we evaluate CART using eight image classification tasks and demonstrate that it significantly outperforms existing model merging approaches, both in the settings with and without test-time adaptation. Our contributions are summarized as follows:
@@
  • We conduct a spectral analysis of the task vectors induced by weight averaging, revealing that the singular vectors effectively isolate core knowledge from interfering noise in each task.
  • Our experiments on eight image classification tasks demonstrate that this simple yet effective approach significantly outperforms prior model merging methods, reducing the performance gap with traditional multi-task learning to within 1–3%.
The remaining sections are organized as follows. We first review model merging approaches including weight averaging and task arithmetic in Section 3. Section 4 introduces intriguing observations of weight averaging along with our analyses, leading to the derivation of our method. Section 5 presents and discusses experimental results on image classification tasks.
@@ 2 Related Work
To overcome the limitations of traditional multi-task learning, various methods have been proposed to construct a single model capable of handling multiple tasks by manipulating model parameters alone [28, 33, 11]. However, due to the non-convexity of loss functions and the complexity of the loss landscape in deep learning tasks, naively summing the parameters of models trained on different tasks often leads to significant performance degradation on each task [17, 23]. Intriguingly, Ilharco et al. [12] demonstrated that arithmetic operations in parameter space using task vectors—differences between the parameters of task-specific fine-tuned models and those of a common pretrained model—consistently influence model behavior. Specifically, adding task vectors enhances performance on multiple tasks (task addition), while subtracting a task vector diminishes performance on the associated task (task negation). This suggests that task vectors capture task-specific knowledge in a way that can be algebraically manipulated. Ortiz-Jimenez et al. [24] analyzed this phenomenon, attributing it to weight disentanglement—where different directions in the weight space of a model independently govern distinct regions in the input space, allowing the model to manipulate these regions separately for specific tasks—and described it as an emergent property inherent to pretrained parameters.
@@
Building upon the findings of Ilharco et al. [12], multi-task models can be constructed without retraining or access to training data by performing a weighted sum of task vectors. However, task interference still leads to reduced performance on individual tasks. To mitigate this issue, recent studies have proposed methods such as Ties-Merging [35], which trims parameters with small magnitudes in the task vectors and resolves sign conflicts among parameters with large magnitudes through majority voting across tasks. Similarly, Yang et al. [36] introduced a method that employs test-time adaptation [31] to dynamically determine the coefficients for the weighted sum of task vectors based on the input data distribution, thereby reducing task interference. Nonetheless, these methods are limited in further reducing task interference because they rely on the fixed definition of task vectors from the pretrained parameters.
@@ 3 Preliminary
We address the problem of model merging, which seeks to combine multiple task-specific models into a single model capable of performing all tasks. Let $\theta_{1},\cdots,\theta_{T}$ denote the parameters of the models trained for $T$ different tasks and $\mathcal{L}_{t}(\theta)$ represent the corresponding loss function for task $t$. Our objective is to merge the parameters to produce $\theta_{*}$ that can perform all tasks, i.e., $\mathcal{L}_{t}(\theta_{*})\approx\mathcal{L}_{t}(\theta_{t})$ for each task $t=1,\dots,T$. By merging models in the parameter space, we need neither an access to training data nor maintaining all individual model parameters. We adopt the standard setting [20, 13, 12] where the models share the same architecture and are fine-tuned from a common initialization $\theta_{0}$, e.g., a pre-trained vision transformer [5, 26].
@@ 4.1 Intriguing Properties of Weight Averaging
Figure 2: An average performance of model merging with reverse-ordered low-rank approximations of task vectors.
(a) Proper-ordered singular vectors
(b) Reverse-ordered singular vectors
Figure 3: Average performances of direct routing with proper- and reverse-ordered low-rank approximation of task vectors.
@@ 4.2 Analysis on Singular Vectors
To understand the intriguing properties of weight averaging discussed in Section 4.1, we propose a conjecture on the intrinsic structures of the task vectors with further analysis. We hypothesize that the interesting behavior of centered task vectors comes from a separation of knowledge contained in their singular vectors. The type of information represented by task vectors can be regarded as either core knowledge that is necessary to solve each task, or noise that comes from the randomness of fine-tuning. Our conjecture is that in the centered task vectors $\bar{\tau}_{t}$, core knowledge is concentrated in top singular vectors, while noise is dispersed (ambient) in lower singular vectors. Conversely, in the original task vectors $\tau_{t}$, we suspect that both core knowledge and noise are spread across all singular vectors.
Such isolation of core knowledge and noise in centered task vectors would allow their low-rank approximations to effectively retain the essential task-specific information with less interference, which explains the ”bell-curve” behavior in Figure 1. When the rank is small, the performance of the centered task arithmetic rapidly increases by continuously absorbing the core knowledge. After absorbing all core knowledge, the remaining noise components cause severe interference in the parameter space, thus the performance drops with larger rank. At the full rank, the core knowledge and noise exactly counteract each other and the merged model reduces to the weight average. On the other hand, this explains why the performance of the original task arithmetic monotonically increases with rank; the core knowledge is distributed across all singular vectors.
To consolidate the conjecture, we perform the low-rank approximation in reverse order, where we preserve lowest $k$ singular vectors for computing $\text{SVD}_{k}$ in Eq. (4). The result of the arithmetic with reverse-ordered rank reduction is shown in Figure 2. We observe an inverted bell-curve of the centered task arithmetic (orange solid line), where the accuracy gradually decreases as the rank increases, followed by a rapid rise back to the weight averaging at high ranks near the full rank. This supports our conjecture, where the gradual performance drop is attributed to the noise from lower singular vectors, and the rapid rise near the full rank is attributed to the core knowledge from top singular vectors. The behavior of the original task arithmetic (blue solid line) is also consistent with our conjecture, where the performance still increases monotonically as rank increases, due to the distributed knowledge.
To further validate the conjecture, we apply direct routing of task vectors with low-rank approximations. In this setting, we maintain each task vector separately rather than merging the task vectors into a single model. Then to perform each task $t$, we route the corresponding task vector with low-rank approximation:
$\mathcal{R}_{k}(t;\theta)=\theta^{l}+\text{SVD}_{k}(\theta_{t}^{l}-\theta^{l}),$ (5)
with either $\theta^{l}=\theta_{\text{avg}}^{l}$ or $\theta^{l}=\theta_{0}^{l}$. Figure 3(a) and  3(b) illustrate the average accuracy of the routed models with proper- and reverse-ordered low-rank approximations, respectively. From Figure 3(a), we observe that the performance using centered task vectors (orange solid line) saturates in lower rank than those using original task vectors (blue solid line). This shows again that the core knowledge is concentrated to top singular vectors of centered task vectors. Similarly, Figure 3(b) supports that their lower singular vectors contain noise, which contributes to the slow increases in performance for most of the ranks.
@@ 4.3 Model Merging with CART
Based on the previous discussions, we propose a training-free model merging approach called Centered Arithmetic with Rank-reduced Task vectors (CART). The method is simple; given $T$ parameters $\theta_{1},\cdots,\theta_{T}$ individually trained for each task $t$, we apply Eq. (4) to obtain merged weight matrices of the model. Compared to the original task arithmetic, CART introduces an additional hyperparameter $k$. However, as we discuss in the experiment (Figure 5(a)), retaining 8% of the rank yields stable performance across different settings.
@@
which we refer as CART++ in our experiments.
@@ 5 Experiments
In this section, we present our experimental results. First, we describe the experimental setup in Section 5.1, followed by a comparison of our method with state-of-the-art model merging techniques in Section 5.2. Finally, we provide an in-depth analysis of our method, exploring its unique behavior, scalability, and sensitivity in Section 5.3.
@@ 5.1 Experimental Setup
Tasks and Models
Following the prior arts in model merging [12, 35, 36], we conducted experiments on eight vision classification tasks spanning diverse domains and varying numbers of labels: Cars [14], DTD [4], EuroSAT [9], GTSRB [29], MNIST [15], RESISC45 [2], SUN397 [34], and SVHN [22]. Additionally, to examine the effects across different model scales, we employed the ViT-B-32 and ViT-L-14 architectures of CLIP [26]. Following the previous settings [12, 35, 36], we exploit task-specific classification heads obtained by the CLIP text encoder, to evaluate the merged model in each task. We report the classification accuracy at each task and their average.
@@
Following Yang et al. [36], we categorize our baselines into several groups. The first group consists of non-merging approaches, including zero-shot application of the pre-trained model (Pretrained), individually fine-tuned models for each task (Individual), and multi-task learning with joint training (Traditional MTL), which respectively represent the lower- and upper-performance bounds for model merging. The second group includes weight averaging methods, such as naive Weight Averaging and its advanced variations, including Fisher Merging [20] and RegMean [13]. The third group includes approaches based on task vectors such as Task Arithmetic [12] and Ties-Merging [35] that aims to reduce task interference. For task vector-based approaches, we additionally consider the method with test-time adaptation that adapts the merging coefficient layer-wise such as AdaMerging [36].
@@
When merging the model, we applied the low-rank approximation of Eq. (4) to only the matrix component of the parameters such as weight matrices in MLP and project matrices in attention layers, while non-matrix components such as biases or ones in normalization layers are set to standard weight averaging. For hyperparameters, we observe that choosing $k=0.08\cdot r$, (i.e., 8% of full rank) and $\lambda=1.0$ works consistently the best across datasets and models.
Table 1: Multi-Task Performance on Eight Vision Tasks with Merged ViT-B-32 with CART.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
Pretrained 62.3 59.7 60.7 45.5 31.4 32.6 48.5 43.8 48.0
Individual 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4 90.5
Traditional MTL 73.9 74.4 93.9 98.2 95.8 98.9 99.5 77.9 88.9
Without Test-Time-Adaptation
Weight Averaging 65.2 63.4 71.5 71.9 64.2 52.8 87.5 50.7 65.9
Fisher Merging [20] 68.6 69.2 70.7 66.4 72.9 51.1 87.9 59.9 68.3
RegMean [13] 65.3 63.5 75.6 78.6 78.1 67.4 93.7 52.0 71.8
Task Arithmetic [12] 55.2 54.9 66.7 78.9 80.2 69.7 97.3 50.4 69.1
Ties-Merging [35] 59.8 58.6 70.7 79.7 86.2 72.1 98.3 54.2 72.4
CART (Ours) 68.1 71.1 84.7 96.0 87.5 89.4 98.9 68.4 83.0
With Test-Time-Adaptation
AdaMerging [36] 64.5 68.1 79.2 93.8 87.0 91.9 97.5 59.1 80.1
AdaMerging++ [36] 66.6 68.3 82.2 94.2 89.6 89.0 98.3 60.6 81.1
CART++ (Ours) 69.5 75.1 89.3 95.7 93.0 96.8 98.9 74.5 86.6
Table 2: Multi-Task Performance on Eight Vision Tasks with Merged ViT-L-14 with CART.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
Pretrained 66.8 77.7 71.0 59.0 58.4 50.5 76.3 55.3 64.5
Individual 82.3 92.4 97.4 100.0 98.1 99.2 99.7 84.1 94.2
Traditional MTL 80.8 90.6 96.3 96.3 97.6 99.1 99.6 84.4 93.5
Without Test-Time-Adaptation
Weight Averaging 72.1 81.6 82.6 91.9 78.2 70.7 97.1 62.8 79.6
Fisher Merging [20] 69.2 88.6 87.5 93.5 80.6 74.8 93.3 70.0 82.2
RegMean [13] 73.3 81.8 86.1 97.0 88.0 84.2 98.5 60.8 83.7
Task Arithmetic [12] 73.9 82.1 86.6 94.1 87.9 86.7 98.9 65.6 84.5
Ties-Merging [35] 76.5 85.0 89.3 95.7 90.3 83.3 99.0 68.6 86.0
CART (Ours) 77.3 89.9 93.5 98.8 94.3 95.9 99.4 77.4 90.8
With Test-Time-Adaptation
AdaMerging [36] 79.0 90.3 90.8 96.2 93.4 98.0 99.0 79.9 90.8
AdaMerging++ [36] 79.4 90.3 91.6 97.4 93.4 97.5 99.0 79.2 91.0
CART++ (Ours) 80.0 92.0 94.0 98.8 96.1 98.5 99.2 82.0 92.6

### v2

@@ Abstract
Model merging aims to build a multi-task learner by combining the parameters of individually fine-tuned models without additional training. While a straightforward approach is to average model parameters across tasks, this often results in suboptimal performance due to interference among parameters across tasks. In this paper, we present intriguing results that weight averaging implicitly induces task vectors centered around the weight averaging itself and that applying a low-rank approximation to these centered task vectors significantly improves merging performance. Our analysis shows that centering the task vectors effectively reduces task interference and most of task-specific knowledge is concentrated in the top singular vectors. Our method demonstrates robust and scalable performance on vision benchmarks across varying numbers of tasks and model sizes. Furthermore, we observe that our approach is applicable to natural language processing tasks with competitive performance.
@@ 1 Introduction
Model merging has emerged as an efficient way to construct multi-task learners [26]. Unlike traditional multi-task learning approaches that directly train a single model on multiple tasks [1], model merging leverages individually fine-tuned models and fuses their parameters to create a model that preserves their original capabilities. This approach eliminates the need to prepare training data or store separate sets of parameters for each task, thereby reducing the costs associated with modern deep neural networks, which often require large amounts of data and numerous parameters. Consequently, model merging has been favored in various applications, including federated learning [37], model compression [14, 47], and continual learning [15, 29].
@@
To mitigate the limitations of the weight averaging, recent approaches have leveraged the task arithmetic framework [17], which allows for the extrapolation of the parameters. In this framework, models are assumed to be fine-tuned from a common initialization, enabling the definition of task vectors as the directions pointing from the initialization to the fine-tuned parameters in the parameter space. Among several arithmetic operations on task vectors, addition with a scaling coefficient has shown to be effective in merging models trained on various tasks. Subsequent approaches have improved this arithmetic by either resolving interference between task vectors [51, 11] or applying test-time adaptation techniques for better scaling coefficient [52].
In this paper, we revisit the weight averaging strategy through the lens of task arithmetic. We begin by formulating weight averaging as a task arithmetic that induces centered task vectors around the weight average itself. We then observe that applying low-rank approximations on the centered task vectors dramatically improves the performance of the merged model produced by the corresponding task arithmetic. When an appropriate rank is chosen, this approach largely outperforms the original task arithmetic and even advanced task vector-based variants. We also note that the optimal rank can be consistently found throughout different experimental settings. To understand this surprisingly high performance, we provide both theoretical insights and empirical evidence based on the spectral analysis of task vectors. With analysis on task interference and spectral analysis of task vectors, we show that centering the task vectors effectively reduces task interference, and most of the task-specific knowledge is concentrated in the top singular vectors. Based on this observation, we propose a novel training-free model merging approach called Centered Arithmetic with Rank-reduced Task vectors (CART), which consistently outperforms existing baselines across various tasks and model architectures. This method can be seamlessly integrated into existing merging approaches based on task vectors, such as those using test-time adaptation. In our experiments, we evaluate CART using various vision and NLP tasks and demonstrate that it significantly outperforms existing model merging approaches, both in the settings with and without test-time adaptation. Our contributions are summarized as follows:
@@
  • We conduct analysis regarding inteference and spectral aspect of task vector, revealing that centering the task vectors reduces task interference.
  • Our extensive experiments on vision and NLP benchmarks with varying numbers of tasks demonstrate that this simple yet effective approach shows competitive performance over prior model merging methods.
@@ 2 Related Work
To overcome the limitations of traditional multi-task learning, various methods have been proposed to construct a single model capable of handling multiple tasks by manipulating model parameters alone [40, 48, 16]. However, due to the non-convexity of loss functions and the complexity of the loss landscape in deep learning tasks, naively summing the parameters of models trained on different tasks often leads to significant performance degradation on each task [25, 33].
Intriguingly, Ilharco et al. [17] demonstrated that arithmetic operations in parameter space using task vectors—differences between the parameters of task-specific fine-tuned models and those of a common pretrained model—consistently influence model behavior. Specifically, adding task vectors enhances performance on multiple tasks (task addition), while subtracting a task vector diminishes performance on the associated task (task negation). This suggests that task vectors capture task-specific knowledge in a way that can be algebraically manipulated. Ortiz-Jimenez et al. [35] analyzed this phenomenon, attributing it to weight disentanglement—where different directions in the weight space of a model independently govern distinct regions in the input space, allowing the model to manipulate these regions separately for specific tasks—and described it as an emergent property inherent to pretrained parameters.
@@
Building upon the findings of Ilharco et al. [17], multi-task models can be constructed without retraining or access to training data by performing a weighted sum of task vectors. However, task interference still leads to reduced performance on individual tasks. To mitigate task interference in multi-task model merging, recent studies have proposed methods that directly operate on task vectors: Ties-Merging [51] employs element-wise manipulations by trimming parameters with small magnitudes and resolving sign conflicts among parameters with large magnitudes via majority voting across tasks; similarly, Consensus Merging [47] leverages task-specific binary masks to retain only those parameters deemed important by multiple tasks, thereby reducing interference-inducing elements. Recently, Yang et al. [52] introduced a method that employs test-time adaptation [46] to dynamically determine the coefficients for the weighted sum of task vectors based on the input data distribution, thereby reducing task interference. Nonetheless, these methods are limited in further reducing task interference because they rely on the fixed definition of task vectors from the pretrained parameters.
@@ 3 Preliminary
We address the problem of model merging, which seeks to combine multiple task-specific models into a single model capable of performing all tasks. Let $\theta_{1},\cdots,\theta_{T}$ denote the parameters of the models trained for $T$ different tasks and $\mathcal{L}_{t}(\theta)$ represent the corresponding loss function for task $t$. Our objective is to merge the parameters to produce $\theta_{*}$ that can perform all tasks, i.e., $\mathcal{L}_{t}(\theta_{*})\approx\mathcal{L}_{t}(\theta_{t})$ for each task $t=1,\dots,T$. By merging models in the parameter space, we need neither an access to training data nor maintaining all individual model parameters. We adopt the standard setting [30, 18, 17] where the models share the same architecture and are fine-tuned from a common initialization $\theta_{0}$, e.g., a pre-trained transformer [8, 38].
@@ 4.1 Intriguing Properties of Weight Averaging
Figure 2: For the ViT-B/32 model with 8 vision tasks, we compute the row space interference $I(k)$ for both centered and original task vectors. The plots illustrate the results for a layer, showing that the centered task vectors consistently exhibit lower $I(k)$ values compared to the original task vectors across all ranks $k$. Layer-wise means and detailed distribution plot for each layer is provided in the Appendix C.1.
Figure 3: For the ViT-B/32 model with 8 vision tasks, the reconstruction error $R(k)$ and row space interference $I(k)$ are computed for the centered task vectors. In the results obtained for a layer, it is observed that as the rank $k$ increases, $R(k)$ exhibits a sharp decline in the low-rank regime. Concurrently, $I(k)$ demonstrates a gradual rise with increasing rank.
@@ 4.2 Analysis on Singular Vectors
To understand the intriguing properties of weight averaging discussed in Section 4.1, we first examine in depth the relationship between interference among task vectors in the perspective of orthogonality. Based on the intuition that task interference is minimized when each task vector is spanned by nullspace of the other task vectors, we define Row space Interference $I(k)$ as follow:
$I(k):=\sum_{i=1}^{T}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{T}\left\|\left(\tilde{\Sigma}_{i}\tilde{V}_{i}\right)^{T}\left(\tilde{\Sigma}_{j}\tilde{V}_{j}\right)\right\|_{F}$ (5)
where $\tilde{V}_{t}$ is the top-$k$ right singular vectors of the task vector $\tau_{t}$, and $\tilde{\Sigma}_{t}$ is a diagonal matrix with corresponding singular values normalized by their $\ell_{2}$ norm. Then, for a single layer, we can derive a simple theorem establishing the relationship between $I$ and task interference as follows:
Theorem 1.
Assume a multi-task model defined as $\theta_{\mathrm{MTL}}=\theta_{0}+\sum_{t=1}^{T}\tau_{t}$ and inputs $x_{t,i}$ to each task $t$ lie close to the subspace spanned by row space of $\tau_{t}$. If we define Task Interference $L:=\sum_{t=1}^{T}\sum_{i=1}^{n}\|\theta_{\mathrm{MTL}}x_{t,i}-(\theta_{0}+\tau_{t})x_{t,i}\|_{2}^{2}$, then following bound holds:
$L\leq\mathcal{O}(I^{2})$ (6)
A more strict definition and the proof are provided in Appendix A. Ideally, achieving $L=0$ implies that the input for each task is transformed exclusively by its own task vector without being influenced by other tasks. Consequently, a smaller value of $I(k)$ implies reduced interference among tasks. In this context, Figure 2 illustrates that, across all ranks, the centered task vectors consistently exhibit lower $I(k)$ compared to the original task vectors. This observation provides insight regarding the superior performance of centered task vectors in Figure 1, since reducing Row Space interference $I$ lowers Task interference by Theorem 1. To further understand the phenomenon that centered task vectors regress toward the average as the rank increases, we define Reconstruction Error $R(k)$ as:
$R(k):=\sum_{i=1}^{T}\left\|\,\theta_{i}-\theta_{avg}-\text{SVD}_{k}(\theta_{i}-\theta_{avg})\,\right\|_{\text{F}}^{2}.$ (7)
Ideally, $R(k)=0$ implies perfect reconstruction of the corresponding task vectors. We examine $R(k)$ and $I(k)$ together in Figure 3, observing that $R(k)$ exhibits a steep decline in the low-rank regime. It indicates that a small number of top singular vectors reconstruct the centered task vectors sufficiently with low task interference. However, beyond this low-rank regime, the improvement in $R(k)$ becomes marginal and the increase in $I(k)$ begins to dominate. Ultimately, this interaction elucidates the regression of the centered task vectors toward the average, explaining the ”bell-curve” shape of Figure 1.
@@ 4.3 Model Merging with CART
Based on the previous discussions, we propose a training-free model merging approach called Centered Arithmetic with Rank-reduced Task vectors (CART). The method is simple; given $T$ parameters $\theta_{1},\cdots,\theta_{T}$ individually trained for each task $t$, we apply Eq. (4) to obtain merged weight matrices of the model. Compared to the original task arithmetic, CART introduces an additional hyperparameter $k$. However, as we discuss in the experiment (Figure 6), retaining 8% of the rank yields stable performance across different settings.
@@
which we refer as Adamerging+CART in our experiments.
@@ 5 Experiments
In this section, we present our experimental results. First, we describe the experimental setup in Section 5.1, followed by a comparison of our method with state-of-the-art model merging techniques in Section 5.2. Finally, we provide an in-depth analysis of our method, exploring its unique behavior, scalability, and sensitivity in Section 5.3. To ensure reproducibility of our results, we release the source code. [^1] https://github.com/JH-GEECS/CART_public
@@ 5.1 Experimental Setup
Computer Vision
To systematically evaluate the scalability and generalization of our model merging approach, we conducted experiments across diverse visual classification tasks with varying domains and label distributions. Initially, we considered an 8-task configuration comprising Cars [21], DTD [4], EuroSAT [13], GTSRB [42], MNIST [23], RESISC45 [2], SUN397 [50], and SVHN [32]. To further examine scalability, we expanded this setting to 14 tasks by incorporating CIFAR100 [22], FER2013 [20], Flowers102 [34], OxfordIIITPet [36], PCAM [43], and STL10 [6]. Subsequently, we extended our evaluation to a 20-task setup by adding EMNIST [7], CIFAR10 [22], Food101 [19], FashionMNIST [49], RenderedSST2 [41], and KMNIST [5] following prior practice [47].
@@
we categorize our baselines into two groups. The first group consists of non-merging approaches, including zero-shot application of the pre-trained model (Pretrained) and individually fine-tuned models for each task (Individual) which respectively represent the lower- and upper-performance bounds for model merging. The second group includes Weight Average and task vector based approaches such as Task Arithmetic [17], Ties-Merging [51] and Consensus Merge [47] that aims to reduce task interference. Also, we additionally consider the method with test-time adaptation that adapts the merging coefficient layer-wise such as AdaMerging [52].
Natural Language Processing
To further assess the generalizability of our method beyond the vision modality, we additionally performed experiments on eight natural language processing (NLP) classification tasks from the GLUE benchmark [45]. Specifically, we merged fully fine-tuned task-specific models built upon the RoBERTa-base backbone [28], preserving the individual task-specific classification heads. This approach enables task-wise evaluation of the merged backbone. We compared our method against task arithmetic-based model merging techniques, including weight averaging, task arithmetic [17], and Ties-Merging [51]. For each of these approaches, we report reproduced results using the hyperparameter settings suggested in their original papers.
@@
When merging the model, we applied the low-rank approximation of Eq. (4) to only the matrix component of the parameters such as weight matrices in MLP and project matrices in attention layers, while non-matrix components such as biases or ones in normalization layers are set to standard weight averaging.
Table 1: Multi-Task Performances on 8, 14 and 20 vision tasks with merged ViT-B/32 and ViT-L/14.
Method ViT-B/32 ViT-L/14
8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
Pretrained $48.0$ $57.1$ $56.0$ $65.0$ $68.4$ $65.4$
Individual $90.5$ $89.6$ $90.5$ $94.2$ $93.3$ $94.1$
Without Test Time Adaptation
Weight Averaging $65.9$ $64.3$ $61.0$ $79.6$ $76.8$ $71.7$
Task Arithmetic [17] $69.1$ $65.4$ $60.5$ $84.5$ $79.6$ $74.2$
Ties-Merging [51] $72.4$ $65.2$ $62.9$ $86.1$ $79.5$ $75.8$
Consensus TA [47] $75.2$ $70.0$ $65.0$ $86.6$ $81.9$ $78.8$
CART (Ours) $\mathbf{84.7}$ $\mathbf{79.5}$ $\mathbf{76.3}$ $\mathbf{92.6}$ $\mathbf{88.7}$ $\mathbf{87.9}$
With Test Time Adaptation
Adamerging+TA [52] $80.1$ $76.7$ $69.2$ $90.8$ $88.0$ $86.8$
Adamerging+CART (Ours) $\mathbf{85.8}$ $\mathbf{82.3}$ $\mathbf{82.7}$ $\mathbf{93.1}$ $\mathbf{90.4}$ $\mathbf{91.3}$
Figure 4: RADAR plot illustrating performance on sets of 8, 14, and 20 vision classification tasks for ViT-B/32, with individual task accuracies normalized by their respective single-task performance. Corresponding results for ViT-L/14 are presented separately in Appendix C.3.